OpenStudy (anonymous):

Prove <1 is congruent to <2

OpenStudy (anonymous):

 Given that \[\frac{ PR }{ MR} = \frac{ QR }{ NR} \]

OpenStudy (michele_laino):

hint: starting from your condition, we can write: \[\begin{gathered} \frac{{MR}}{{PR}} = \frac{{NR}}{{QR}} \hfill \\ \left( {MR - PR} \right):PR = \left( {NR - QR} \right):NR \hfill \\ MP:PR = NQ:NR \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

Now we can apply the theorem of Talete. What can we conclude?